# On the complexity of 2-bridge link complements

### v1

@@ 5.2 Angle structures on blocks
where the deficits $x_{i}$ are as in Equation 5.16. Assigning $\Theta^{(2)}=(\pi/2,\pi/4,\pi/4)$ and $\Theta^{(3)}=(\pi/4,\pi/4,\pi/2)$ gives $\Delta_{1}$ shape (VII) as above if $y=\pi/4$. Note that $\Delta_{4}$ must have either shape (III) or shape (VI) and $y$ must correspond to the diagonal edge class. In either case, we have $\theta_{3(4)}=\pi/4$ and thus the shape assignment is compatible.
@@
Without loss of generality, assume the last letter of $\Omega$ is $R$. Suppose that $k=2$ and the (B2) block is given by $R^{2}L^{2}$. The block contains the layers above with $\Delta_{\Sigma-2}$ and $\Delta_{\Sigma-3}$ now added. The angle equations of this block are now,
@@
Increasing to $k=3$ the (B2) block is now given by $L^{2}R^{2}L^{2}$. We can observe that the last three angle equations above remain the same. The three boundary angle equations are the same as above, however the angles assigned to vertical and horizontal edges are swapped. From this we assign shape (IX) to $\Delta_{\Sigma-5}$ as above, but swapping the first two angles and shape (V) to layers $\Delta_{\Sigma-2}$, $\Delta_{\Sigma-1}$ and $\Delta_{\Sigma}$ exactly as above. The angle equations involving layers $\Delta_{\Sigma-4}$ and $\Delta_{\Sigma-3}$ are,
@@
An analogous argument shows that this assignment of shapes is valid if the (B2) block ends on $R^{2}$ or starts on $L^{2}$ with the first two angles of each shape vector swapped. ∎
@@ 5.3 Complexity bounds
$\frac{\mathcal{V}(\Theta_{1}^{(B2)})}{2(2k_{1}+1)v_{3}}=\frac{2\left(v_{VII}+v_{I}+v_{VI}+2(k_{0}-1)v_{III}\right)}{2(2k_{1}+1)v_{3}}=\frac{3.6096k_{1}+1.405}{2(2k_{1}+1)}.$
@@
$\frac{\mathcal{V}(\Theta_{1}^{(B3)})}{2(3m_{1}+2)v_{3}}=\frac{2\left(v_{V}+v_{I}+v_{II}+v_{IV}+v_{VI}+2(m_{0}-1)v_{III}+(m_{1}-1)v_{VIII}\right)}{2(3m_{0}+2)v_{3}}=\frac{4.9632m_{1}+3.9264}{2(3m_{1}+2)}.$
@@
This evaluates to $0.8381$ for $k=2$ and increases with $k$ and satisfies our desired result. However, when $k=1$ this evaluates to $0.7952$. In this case we assign a different angle structure to the triangulation. We solve the angle equations, obtained from Figures 4 and 5, by assigning
@@
It follows that the Sakuma-Weeks triangulation of a $2$-bridge link associated to the work $\Omega=RL^{a_{1}}\cdots(L^{a_{n}}R\mid R^{a_{n}}L)$ with $a_{i}\in\{1,2\}$ for all $1\leq i\leq n$ satisfies $0.8|\mathcal{T}|\leq c(M)\leq|\mathcal{T}|$. ∎
@@
It follows from the proof of Theorem 25 that the only case we need to consider is when $\Omega$ decomposes as a single (B1) block of length $13C+$ 1. Let $\Phi^{*}$ denote the associated angle structure. From LABEL:eq:SingleB3 we compute the largest deficit obtained from $C$ of the $a_{i}=2$ as

### v2

@@ 5.2 Angle structures on blocks
where the deficits $x_{i}$ are as in Equation 5.16. Assigning $\Theta^{(2)}=(\pi/2,\pi/4,\pi/4)$ and $\Theta^{(3)}=(\pi/4,\pi/4,\pi/2)$ gives $\Delta_{1}$ shape (VII) as above if $z=\pi/4$. Note that $\Delta_{4}$ must have either shape (III) or shape (VI) and $z$ must correspond to the diagonal edge class. In either case, we have $\theta_{3(4)}=\pi/4$ and thus the shape assignment is compatible.
@@
Assume the last letter of $\Omega$ is $R$. Suppose that $k=2$ and the (B2) block is given by $R^{2}L^{2}$. The block contains the layers $\Delta_{\Sigma-3}$ to $\Delta_{\Sigma}$, inclusive. The angle equations of this block are now,
@@
Increasing to $k=3$ the (B2) block is now given by $L^{2}R^{2}L^{2}$. We can observe that the last three angle equations above remain the same. The three boundary angle equations are the same as above, however the angles assigned to vertical and horizontal edges are swapped. From this we assign shape (IX) to $\Delta_{\Sigma-5}$ as was done for $\Delta_{\Sigma-3}$ above, but swapping the first two angles and shape (V) to layers $\Delta_{\Sigma-2}$, $\Delta_{\Sigma-1}$ and $\Delta_{\Sigma}$ exactly as above. The angle equations involving layers $\Delta_{\Sigma-4}$ and $\Delta_{\Sigma-3}$ are,
@@
An analogous argument shows that this assignment of shapes is valid if the (B2) block ends on $R^{2}$ (giving the last letter of $\Omega$ to be $L$) or starts on $L^{2}$ with the first two angles of each shape vector swapped. ∎
@@ 5.3 Complexity bounds
$\frac{\mathcal{V}(\Theta_{1}^{(B2)})}{2(2k_{1}+1)v_{3}}=\frac{2\left(v_{VII}+v_{I}+v_{VI}+2(k_{1}-1)v_{III}\right)}{2(2k_{1}+1)v_{3}}=\frac{3.6096k_{1}+1.405}{2(2k_{1}+1)}.$
@@
$\frac{\mathcal{V}(\Theta_{1}^{(B3)})}{2(3m_{1}+2)v_{3}}=\frac{2\left(v_{V}+v_{I}+v_{II}+v_{IV}+v_{VI}+2(m_{1}-1)v_{III}+(m_{1}-1)v_{VIII}\right)}{2(3m_{1}+2)v_{3}}=\frac{4.9632m_{1}+3.9264}{2(3m_{1}+2)}.$
@@
This evaluates to $0.8381$ for $k=2$ and increases with $k$ and satisfies our desired result. However, when $k=1$ this evaluates to $0.7952$. In this case we assign a different angle structure to the triangulation. We solve the angle equations, obtained from Figures 4 and 5(a), by assigning
@@
It follows that the Sakuma-Weeks triangulation of a $2$-bridge link associated to the word $\Omega=RL^{a_{1}}\cdots(L^{a_{n}}R\mid R^{a_{n}}L)$ with $a_{i}\in\{1,2\}$ for all $1\leq i\leq n$ satisfies $0.8|\mathcal{T}|\leq c(M)\leq|\mathcal{T}|$. ∎
@@
It follows from the proof of Theorem 25 that the only case we need to consider is when $\Omega$ decomposes as a single (B1) block of length $13C+$ 1. Let $\Phi^{*}$ denote the associated angle structure. The largest deficit is obtained when $\Omega$ decomposes as a single (B3) block of length $13C+$ 1 consisting of $C$ (B2) blocks of length $1$. Let $\Theta^{*}$ denote the associated angle structure. We compute the volume functional for $\Theta^{*}$ as
$\mathcal{V}(\Theta^{*})=2\left(v_{V}+v_{I}+v_{VI}+(C-1)(2v_{III}+v_{VIII})+v_{III}+v_{VIII}\right)\approx(4.9632C+3.3930)v_{3}.$
The largest deficit obtained from $\Omega\in W_{C}$ is then